OpenStudy (anonymous):

Solve using the subsitution method. 4x - 8y = 48 and 11x + 3y = -105.5

OpenStudy (anonymous):

y = -9.5 and x = 31

OpenStudy (anonymous):

Can you please show me how you got that answer ?

OpenStudy (anonymous):

Sure! Pick either equation 1 or 2 (as I've decided to name them, they are numbered in order of appearance). I chose to solve equation 1 for x. This gave me: x = 12 + 2y. I then plugged that value for x into equation 2 such that: 11x + 3y = -105.5 becomes 11(12 +2y) + 3y = -105.5. Solve for y and you get -9.5.

OpenStudy (anonymous):

Now if you take that value for y and plug it into equation 1, solving for x, you get 31. Make sense?

OpenStudy (anonymous):

Yes thank u

OpenStudy (anonymous):

Wait - there is an error in my math. X should equal -7.